OpenStudy (anonymous):

can you give me the steps to help solve w+3/4=1/8 in simplest form

OpenStudy (anonymous):

common denominator/giraffe looks to have the answer

OpenStudy (anonymous):

wouldn't you multiply the reciprocal from both sides. Then you would get what w equals? Sorry if this is wrong, it's just a guess.:)

OpenStudy (anonymous):

reciprocal of 3/4 i mean..

OpenStudy (campbell_st):

which equation is it \[\frac{x + 3}{4} = \frac{1}{8}\] or \[x + \frac{3}{4} = \frac{1}{8}\]